### v1

# A Posterior Sampling Approach to Inferring Indoor Floorplans from Walk Trajectories

Abstract
Neural Radiance Fields (NeRFs) have been remarkably successful at synthesizing novel views of 3D scenes by optimizing a volumetric scene function. This scene function models how optical rays bring color information from a 3D object to the camera pixels. Radio frequency (RF) or audio signals can also be viewed as a vehicle for delivering information about the environment to a sensor. However, unlike camera pixels, an RF/audio sensor receives a mixture of signals that contain many environmental reflections (also called “multipath”). Is it still possible to infer the environment using such multipath signals? We show that with redesign, NeRFs can be taught to learn from multipath signals, and thereby “see” the environment. As a grounding application, we aim to infer the indoor floorplan of a home from sparse WiFi measurements made at multiple locations inside the home. Although a difficult inverse problem, our implicitly learnt floorplans look promising, and enables forward applications, such as indoor signal prediction and basic ray tracing.
1 Introduction
NeRFs mildenhall2021nerf ; nerfInvProb ; mipnerf ; pixelnerf have delivered impressive results in solving inverse problems, resulting in 3D scene rendering. While NeRFs have mostly used pictures (from cameras or LIDARs) to infer a 3D scene, we ask if the core ideas can generalize to the case of wireless signals (such as RF or audio). Unlike camera pixels that receive line-of-sight (LoS) rays, a wireless receiver (e.g., a WiFi antenna on a smartphone) would receive a mixture of LoS and many reflections, called multipath. If the receiver moves, it receives a sequence of $N$ measurements. Using these $N$ wireless measurements, is it possible to learn a representation of the scene, such as the floorplan of the user’s home? An optical NeRF understandably fails since it is not equipped to handle multipath. This paper is focused on redesigning NeRFs so they can learn to image the environment, thereby solving the inverse problem from ambient wireless signals.
A growing body of work specnerf ; winert ; zhao2023nerf2 ; newrf ; luo2022learning ; AV-NeRF is investigating connections between NeRFs and wireless. While none have concentrated on imaging, NeRF2 zhao2023nerf2 and NeWRF newrf have augmented NeRFs to correctly synthesize WiFi signals at different locations inside an indoor space. However, correct synthesis is possible without necessarily learning the correct signal propagation models. We find that NeRFs with adequate model complexity can overfit a function to correctly predict signals at test locations, but this function does not embed the true behavior of multipath signal propagation. We re-design the NeRF’s objective function so that it learns the environment through line-of-sight (LoS) paths and reflections. This teaches the NeRF an implicit representation of the scene, which can then be utilized for various forward tasks, including WiFi signal prediction and ray tracing.
In our model, EchoNeRF, each voxel is parameterized by its opacity $\delta\in[0,1]$ and orientation $\omega\in[-\pi,\pi]$. When trained perfectly, free-space air voxels should be transparent ($\delta=0$), wall voxels should be opaque ($\delta=1$), and each opaque voxel’s orientation should match its wall’s orientation. As measurements, we use the received signal power. Thus, the input to our EchoNeRF model is the transmitter ($Tx$) location, a sequence of known receiver ($Rx$) locations, and the signal power measured at each $Rx$ location. The output of EchoNeRF is an (implicitly learnt) floorplan of the indoor space. We expect to visualize the floorplan by plotting the learnt voxel opacities.
Figure 1: LoS and correct multipath reflections.
Figure 2: (a) Fitting signal power using virtual $Tx$s. (b) Ideally, virtual $Tx$s should be located on wall surfaces. (c) Sparse virtual $Tx$s learnt by NeWRF shown in black/gray dots. (d) Dense virtual $Tx$s learnt by NeRF2. Neither correspond to the true (red) floorplan.
Learning floorplans requires modeling the correct reflections (see Fig. 2) since these reflections help reveal where the walls are. However, without knowledge of the walls, the reflections are difficult to model, leading to a type of chicken and egg problem. Additionally, the number of wireless measurements is relatively sparse compared to the number of pixels measured in image-trained NeRFs. Finally, measured signals will have “blind spots”, meaning that rays that bounced off certain regions of the walls may not have arrived at any of the $Rx$ locations. This leaves gaps or holes in the floorplan and NeRF’s interpolation through these gaps will produce error or blur.
EchoNeRF approaches this problem by modeling the received signal power as a combination of the LoS power and the power from first order reflections. The LoS model is inherited from classical NeRFs. The main departure from past work is in modeling the reflections. Since opaque voxels are unknown during training, the reflection surfaces are not known; hence, the reflection power at the $Rx$ is modeled as an aggregate over all plausible reflections. Given the planar structure of walls, the plausible set of reflections can be heavily pruned to reduce the optimization complexity. Reflections aggregated over this plausible set models the total (LoS + reflection) power at a receiver $Rx$.
EchoNeRF trains to minimize the loss between the modeled and measured power across all $Rx$ locations, and in the process, learns the voxel’s opacities that best explain the measured dataset. Some regularization is necessary to cope with sparse measurements and to ensure smoothness of walls. Lastly, to handle some gradient imbalance issues, EchoNeRF freezes the LoS model once it converges, and uses this intermediate state to partly supervise the reflection model.
To evaluate EchoNeRF, we train on $2.4$ GHz WiFi signals from NVIDIA’s Sionna simulator hoydis2023sionna , with floorplans from the Zillow’s Indoor Dataset (ZIND) cruz2021zillow . Results show consistent improvement over baselines in terms of the estimated floorplan’s IoU and F1 score. Qualitative results show visually legible floorplans without any post-processing. Applying forward functions on the floorplan, EchoNeRF can predict the received signal power for new $\langle Tx,Rx\rangle$ locations (outperforming existing baselines). Lastly, basic ray tracing explains the predictions, offering interpretability to its results.
2 Related Work and Research Scope
Wireless (WiFi) channel prediction using NeRFs. NeWRFnewrf and NeRF2zhao2023nerf2 are recent papers that have used NeRFs to predict the wireless channel impulse response (CIR) Tse_Viswanath_2005 at unknown locations inside a room. Drawing a parallel to optical NeRFs, a voxel’s color in optics becomes a voxel’s transmit power in wireless. The voxel’s density in optics remains the same in wireless, modeling how that voxel attenuates signals passing through it. NeRF2 and NeWRF assign transmit power and attenuation to each voxel such that they best explain the measured CIR. The authors explain that voxels assigned non-zero transmit power will be called virtual transmitters; these voxels represent the reflection points on the walls. However, many assignments are possible that fit the CIR training data, especially when the data is sparse. Fig. 2(a,b) illustrates 2 possible assignments. While the predicted CIRs could achieve low error for all such assignments, only one of the assignments will model the true reflections, forcing the virtual transmitters to be located on the wall surfaces (Fig. 2(b)). We have plotted NeRF2 and NeWRF’s assignment of voxel densities (see Fig. 2(c,d)) to confirm that the high accuracy in CIR prediction is not an outcome of correctly learning the wall layout. Our goal is to repair this important issue, i.e., assign voxel densities that obey the basic physics of wall reflections. Correct voxel assignment leads to the correct layouts, which then makes the (forward) CIR prediction easy.
Neural radiance fields for audio. Another active line of research focuses on predicting room impulse response (RIR) for audio luo2022learning ; INRAS ; AV-NeRF ; brunetto2024neraf . Neural Acoustic Field (NAF) luo2022learning extended the classical NeRF to train on RIR measurements in a room and predict the RIR (magnitude and phase) at new $\langle Tx,Rx\rangle$ locations. NAF identified the possibility of overfitting to the RIR and proposed to learn, jointly, the local geometric features of the environment (as spatial latents) and the NAF parameters. The spatial latents embed floorplan information but a decoder needs to be trained using the partial floorplan data. EchoNeRF requires no floorplan supervision, and secondly, relies entirely on signal power (less informative than RIR) to estimate the floorplan.
Follow up work are embracing more information about the surroundings (pictures NACF ; INRAS ; AV-NeRF ; fewshotvisual , LIDAR scans oufqir2020arkit , meshes and optical NeRFs brunetto2024neraf ), to boost RIR accuracy. Results are steadily improving, however, this sequence of ideas is unaligned with solving the core inverse problem. Our goal is to first invert the signal power to a floorplan, which can then enable CIR/RIR predictions.
Modeling reflections in optical NeRFs Optical NeRFs have tackled reflections refneus ; nerfds ; flipnerf for synthesizing glossy surfaces and mirrors, and for re-lighting rudnev2022nerf ; relight . NeRFRen nerfren proposes to decompose a viewed image into a transmitted and a reflected component. Ref-NeRF verbin2022refnerf also focuses on reflections through a similar decomposition of the transmitted and reflected color, however, the reflected color is modeled as a function of the viewing angle and the surface-normal, resulting in accurate models of specular reflection. Several other papers have developed similar ideas boss2021nerd ; zhang2021nerfactor and the core insight centers around solving a two-component decomposition problem. EchoNeRF faces the challenge of not knowing the number of rays adding up from all possible directions in the environment. Hence, EchoNeRF must solve a many-component decomposition problem by leveraging the physics of multipath signal propagation.
3 EchoNeRF Model
Setup and Overview. At a $Rx$ location, we model the received signal power $\psi$ as
$\psi=\psi_{LoS}+\psi_{ref_{1}}+\cdots+\psi_{ref_{n}}$
where $\psi_{LoS}$ is the power from the direct line-of-sight (LoS) path, and $\psi_{ref_{k}}$ is the aggregate power from all $k^{th}$ order reflections (i.e., all signal paths that underwent exactly $k$ reflections before arriving at the $Rx$)  [^1] This model is an approximation since it relies on averaging measurements as explained in Appendix B. However, for this sensing application involving moving receivers, and when using wideband-multipath systems such as WiFi, this approximation holds well. We assume $M$ fixed transmitters and move the $Rx$ to $N$ known locations and measure $\psi$ at each of them. EchoNeRF accepts $M\times N$ measurements as input and outputs the 2D floor-plan $F$, a binary matrix of size $L\times L$, where $L$ denotes the maximum floorplan length.
We train EchoNeRF on the measured data using our proposed objective function. This function only models the LoS and the first order reflections. We disregard the higher orders since they are very complex to model and contribute, on average, $<6\%$ of the total power (see statistics in Appendix C). The NeRF model we use is a remarkably simple MLP designed to predict the density $\delta\in[0,1]$ and orientation $\omega\in[-\pi,\pi]$ of a specified voxel in the indoor scene. The orientation aids in modeling reflections. The proposed objective function – parameterized by voxel attributes $\langle\delta,\omega\rangle$ and the $\langle Tx,Rx\rangle$ locations – models an approximation of the received power $\psi$ at that $Rx$ location. Minimizing $L_{2}$ loss of this power across all $Rx$ locations trains the MLP. Plotting out all the voxel densities in 2D gives us the estimated floorplan $F$.
3.1 The LoS Model
Friss’ equation antennaBook from electromagnetics models the free-space received power as $P_{r}=\frac{K}{d^{2}}$ where $d$ is the distance of signal propagation, and $K$ is a product of transmit-power, wavelength, and antenna-related constants antennaBook . We model this free-space (LoS) behavior in the NeRF framework through the following equation.
$\psi_{LoS}=K\frac{\displaystyle\prod_{\{i|v_{i}\in{LoS}\}}(1-\delta_{i})}{d^{2}}$ (1)
where $K$ can be empirically measured, and $d$ is the known distance between the $\langle Tx,Rx\rangle.$ The numerator includes the product of voxel densities over all voxels along the LoS ray from $Tx$ to $Rx$ (with an abuse of notion, we write this as $v_{i}\in{LoS}$). This models occlusions. When the LoS path is completely free of any occlusions (i.e., $\delta_{i}=0,\forall i$ where $\{i|v_{i}\in{LoS}\}$), we expect the received power to only be attenuated by the pathloss factor $d^{2}$ (in the denominator). Eq. 1 has a slight difference to classical NeRF’s volumetric scene function. In our case, voxels along the ray do not contribute to the received power (whereas in NeRF, each voxel’s color is aggregated to model the final pixel color at the image). In other words, we have modeled a single transmitter in Eq. 1.
3.2 The Reflection Model
To model reflections, consider a voxel $v_{j}$. Whether $v_{j}$ reflects a ray from the $Tx$ towards $Rx$ depends on (1) $v_{j}$’s density $\delta_{j}$ and orientation $\omega_{j}$, (2) the $\langle Tx,Rx\rangle$ locations, and (4) whether the path from $Tx$ to $v_{j}$, and from $v_{j}$ to $Rx$ are both occlusion-free. Parameterized by these, Eq. 2 models $\psi_{ref}(v_{j})$, which is the received power at $Rx$ due to the signal that reflected off voxel $v_{j}$.
$\psi_{ref}(v_{j})=\delta_{j}f(\theta,\beta)\frac{\displaystyle\prod_{k\in\{Rx:v_{j}\}}(1-\delta_{k})\prod_{l\in\{v_{j}:Tx\}}(1-\delta_{l})}{\left(d_{Tx:v_{j}}+d_{v_{j}:Rx}\right)^{2}}$ (2)
Let us explain this equation briefly. The leading $\delta_{j}$ ensures that voxel $v_{j}$ is not a reflector when $\delta_{j}=0$. The $f(\theta,\beta)$ term models the wave-surface interactions, i.e., how signals get attenuated as a function of the incident angle $\theta$ and how signals scatter as a function of the offset angle $\beta$ (which is the angle between the reflected ray and the direction of the $Rx$ from $v_{j}$). The next two product terms ensure that for the $Rx$ to receive this reflection, the voxels along the $2$ segments ($Tx$ to $v_{j}$ and $v_{j}$ to $Rx$) must be non-opaque; if any $\delta_{k}$ or $\delta_{l}$ equals $1$, that reflection path is blocked, producing no power contribution via this voxel $v_{j}$ to the receiver $Rx$. Finally, the denominator is the squared distance from $Tx$ to $v_{j}$, and from $v_{j}$ to $Rx$, modeling signal attenuation.
To compute the full reflection power, the natural question is: which voxels are contributing to the received power? Geometrically, any opaque voxel can be a plausible reflection point between any $\langle Tx,Rx\rangle$ pair. This is because, for triangle formed by $Tx$, $v_{j}$, and $Rx$, the voxel orientation $\omega_{j}$ can be assigned a direction that bisects the angle at $v_{j}$. For this $\omega_{j}$, the reflected ray will perfectly arrive at the $Rx$. Thus, without the knowledge of orientation and density, the total first order reflection power at the $Rx$ should be modeled as the sum of reflections on all voxels. This makes the optimization problem excessively under-determined.
We address this by modeling $\omega_{j}$ as a discrete value, i.e., $\omega_{j}\in\{1,2\dots K_{\omega}\}$. Larger $K_{\omega}$ is needed when the environment has complex surface orientations; however, most floorplans exhibit perpendicular walls manhattan1 ; manhattan2 ; manhattan3 and $K_{\omega}=4$ is adequate. Once $\omega_{j}$ becomes discrete, the voxels that can produce plausible reflections become far fewer – we call this the “plausible set” $\mathcal{V}$. Fig.  4 visualizes $\mathcal{V}$ and shows $3$ out of many plausible reflections from voxels with $\omega_{j}=45^{\circ}$. Eq. 3 sums up the power from all reflections that occur on the plausible set:
$\psi_{ref_{1}}=\displaystyle\sum_{\{j|v_{j}\in\mathcal{V}\}}\psi_{ref}({v_{j}})$ (3)
Thus, the final modeled power at a specific $Rx$ location becomes $\tilde{\psi}=\psi_{LoS}+\psi_{ref_{1}}$.
3.3 Gradient Issues during Training
Figure 3: Colored stripes define the manifold from which reflections are plausible between $\langle Tx,Rx\rangle$. Voxels located on the manifold form the plausible set $\mathcal{V}$. Dashed lines show plausible reflections.
Figure 4: EchoNeRF’s two-stage training approach: In Stage 1, the LoS model is trained using known $Rx$ locations and signal power. This provides a warm-start to the reflection model in Stage 2 which refines the learned voxel densities and orientation.
Training against a $L_{2}$ loss, $\mathcal{L}=\|\tilde{\psi}-\psi^{*}\|^{2}_{2}$  [^2] Here, $\psi^{*}$ denotes the ground truth signal power, did not generate legible floorplans. We found that the $\psi_{LoS}$ dominated the loss term, drowning the reflection model’s influence on learning. At a high level, the gradient of the LoS model (Eq. 1) w.r.t. $\delta_{i}$ has fewer terms in the numerator’s product, and a smaller $d^{2}$ in the denominator. The reflection model’s gradient w.r.t. $\delta_{i}$ has many more terms since the reflection path is much longer; the denominator is also larger. Since $(1-\delta_{j})\leq 1$, their products force the gradient to decrease geometrically with more terms, causing the reflection gradient to be much smaller compared to LoS. We formalize this explanation below by considering the LoS and reflection losses individually [^3] For the ease of explanation, we use $\psi^{*}_{LoS}$ and $\psi^{*}_{ref}$ to denote the ground truth LoS and reflection powers, respectively. We do not need to know these terms in practice..
$\mathcal{L}_{LoS}$ $=\left\|\tilde{\psi}_{LoS}-\psi^{*}_{LoS}\right\|^{2}_{2},\quad\mathcal{L}_{ref}=\left\|\tilde{\psi}_{ref}-\psi^{*}_{ref}\right\|^{2}_{2}$
Consider the gradient of $\mathcal{L}_{LoS}$ w.r.t the density of $v_{i}$.
$\nabla_{\delta_{i}}\mathcal{L}_{LoS}$ $=2(\tilde{\psi}_{LoS}-\psi^{*}_{LoS})\sum_{\{n|i\in LoS^{(n)}\}}\nabla_{\delta_{i}}\psi^{(n)}_{LoS}$
$\text{where}~{}~{}\nabla_{\delta_{i}}\psi^{(n)}_{LoS}$ $=-\frac{K}{{d^{(n)}}^{2}}\prod_{\begin{subarray}{c}j\in LoS^{(n)}\\ j\neq i\end{subarray}}(1-\delta_{j})$ (4)
where $LoS^{(n)}$ is the $n$-th LoS path passing through voxel $v_{i}$.
The gradient of $\mathcal{L}_{ref}$ has a nearly identical expression with the only difference being many more product terms and that $\nabla_{\delta_{i}}\psi_{ref}^{(n)}$ depends on where $v_{i}$ is present in the $n$-th set of reflection path voxels (denoted by $Ref^{(n)}$). For example,
$\nabla_{\delta_{i}}\psi_{ref}^{(n),Tx}$ $=-C\prod_{\begin{subarray}{c}j\in Ref^{(n)}_{Tx:v}\\ j\neq i\end{subarray}}(1-\delta_{j})\prod_{k\in Ref^{(n)}_{v:Rx}}(1-\delta_{k})$ (5)
$C$ $=\frac{\delta^{(n)}f^{(n)}(\theta,\beta)}{(d^{(n)}_{Tx:v}+d^{(n)}_{v:Rx})^{2}}$
here $\nabla_{\delta_{i}}\psi_{ref}^{(n),Tx}$ denotes the gradient when $v_{i}$ is between Tx to the reflection point $v$.
Finally, since the modeled power approximates the measured power, the residual error will remain non-zero even if the floorplan is accurately learnt. As a result, the optimization is biased towards voxels of higher gradients, i.e., voxels on the LoS path, suppressing the importance of reflections. To address this, we train EchoNeRF in $2$ stages.
3.4 Multi-stage Training
Stage 1: We first use the LoS model against the measured ground truth power $\psi^{*}$ (see Fig. 4). This converges quickly because the network easily learns the transparent voxels ($\delta$=00) located along LoS paths. For LoS paths that are occluded, the network incorrectly learns excessive opaque voxels between the $\langle Tx,Rx\rangle$, but this does not affect the LoS error since the path is anyway occluded. Hence, the outcome is a crude floorplan but a near-perfect LoS power estimate $\tilde{\psi}_{LoS}$. We utilize this $\tilde{\psi}_{LoS}$ in stage $2$ (discussed soon).
As Stage $1$ training progresses, some opaque voxels emerge, offering crude contours of some walls. We estimate a voxel’s spatial gradient, $\nabla\delta_{i}$, and use it to supervise the orientation $\omega_{i}$ of that voxel. The intuition is that a voxel’s orientation – needed to model reflections in Stage $2$ – is essentially determined by the local surface around that voxel. The gradient $\nabla\delta_{i}$ offers an opportunity for weak supervision. Thus, the loss for Stage $1$ is:
$\mathcal{L}_{stage1}$ $=\left\|\psi^{*}-\psi_{LoS}\right\|^{2}_{2}+\mathcal{L}^{+}$ (6)
$\text{where}~{}~{}\mathcal{L}^{+}$ $=\lambda_{1}\underset{\forall j}{\sum}\left\|\nabla\delta_{j}-\omega_{j}\right\|_{2}^{2}+\lambda_{2}\mathcal{L}_{reg}$ (7)
with $\lambda_{1},\lambda_{2}>0$ being tunable hyperparameters. The regularization term $\mathcal{L}_{reg}$ will be discussed soon. Finally, the near-perfect estimate of LoS power, denoted $\tilde{\psi}_{LoS}$, is also carried over to Stage $2$ to ensure the reflection model is penalized when it veers away from this LoS estimate.
Stage 2 focuses on training the reflection model using the following loss function.
$\mathcal{L}_{stage2}=\left\|\tilde{\psi}_{LoS}-\psi_{LoS}\right\|^{2}_{2}+\left\|\psi^{*}-\tilde{\psi}_{LoS}-\psi_{ref_{1}}\right\|^{2}_{2}+\mathcal{L^{+}}$
The first term in the RHS ensures that the Stage $1$’s LoS estimate is honored in Stage $2$. The second term subtracts Stage $1$’s LoS power from the measured power, ($\psi^{*}-\tilde{\psi}_{LoS}$); this models the total power only due to reflections. Our (first order) reflection model $\psi_{ref_{1}}$ is trained to match this aggregate power ($L_{2}$ loss). The supervision on orientation and the regularization terms are the same as in Stage $1$.
$\blacksquare$ Regularization: Floorplans demonstrate significant local similarity in orientation, hence we penalize differences in orientation among neighbors, using a regularization (Eq. 8) similar to Total Variation TV_reg . This can be achieved without additional computational cost to the neural network by directly utilizing voxel orientations obtained from each ray.
$\mathcal{L}_{\text{reg}}=\frac{1}{n_{v}(n_{r}-1)}\sum_{n=1}^{n_{v}}\sum_{i=1}^{n_{r}-1}\left\|{\omega}_{n,i+1}-{\omega}_{n,i}\right\|^{2}_{2}$ (8)
Here $n_{v}$ is the number of voxels queried from the plausible set $\mathcal{V}$ and $n_{r}$ is the number of voxels along the each ray.
4 Experiments
Floor plan and Wireless Simulation Dataset. Floorplans are drawn from the Zillow Indoor Dataset  cruz2021zillow . In each floorplan, we use the $A^{*}$ algorithm  astar to generate a walking trajectory that traverses all rooms. We use the NVIDIA Sionna RT hoydis2022sionna ; hoydis2023sionna – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)). We randomly place $M$ Txs, one in each room, denoted as $T_{m}$. To simulate omnidirectional transmissions at $2.4$GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan. For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$. Sionna accounts for specular reflections and refraction when these rays interact with walls in the specified floor plan; we use the default materials for the walls. As with most WiFi simulators, Sionna does not model signal penetration through walls – this means that a Tx and Rx located on opposite sides of a wall will not receive any RSSI. Overall, we gather $M\times N$ RSSI measurements $(T_{m},R_{n},\psi_{m,n})$ that serve as input to EchoNeRF.
Baselines used for comparison are:
  1. 1.
    NeRF2 zhao2023nerf2 : Models WiFi reflections via virtual transmitters to predict channel impulse response (CIR).
  2. 2.
    Heatmap Segmentation liu2012review : Interpolates CIR across the whole floorplan and applies an image segmentation algorithm (on the interpolated RSSI heatmap) to isolate each room. Essentially, the algorithm identifies the contours of sharp RSSI change since such contours are likely to correspond to walls. Implementation details are included in the Appendix E.4.
  3. 3.
    MLP: Trains an MLP network to directly estimate the RSSI based on Tx and Rx locations.
  4. 4.
    EchoNeRF_LoS: Reports EchoNeRF’s result considering only LoS path (ablation study).
Metrics. We evaluate using $3$ metrics:
(A) Wall Intersection over Union (Wall_IoU): This metric measures the degree to which the predicted walls and the true walls superimpose over each other in the 2D floorplan. The following equation defines the metric:
[…]
where $WP$ denotes the set of predicted wall pixels and $WP^{*}$ denotes the true wall pixels. This is a harsh metric given wall pixels are a small fraction of the total floorplan; if a predicted wall is even offset by one pixel from the true wall, the Wall_IoU drops significantly. IoU iou has often been defined in terms of room pixels (instead of wall pixels); this is an overestimate in our opinion, since predicting even an empty floorplan results in an impressively high IoU.
(B) F1 score metrics : Defined as $F1=\frac{2\times P\times R}{P+R}$, where $P$ is the precision and $R$ is the recall of the bitmap. $P$ and $R$ are defined based on wall pixels, similar as above.
(C) RSSI Prediction Error (RPE): We split all $Rx$ locations into a training and test set. RPE reports the average median RSSI error over all the test locations across floorplans.
4.1 Overall Summarized Results
Table 1 reports comparative results between EchoNeRF and baselines, averaged over $20$ different experiments, using all $3$ metrics. The number of measurements are sparse ($N=2000$ and $N=1000$), given that apartment sizes in our dataset are more than $250,000$ pixels. Mean and standard deviation are reported in the table. EchoNeRF outperforms all models in terms of Wall_IoU and F1 Score. Compared to EchoNeRF_LoS, EchoNeRF demonstrates visible improvements, highlighting the advantage of modeling reflections. The absolute Wall_IoU values are understandably low because the metric penalizes small errors.
NeRF2 is unable to predict the floor plan (opaque voxels) well and is only able to achieve better RPE than EchoNeRF_LoS. EchoNeRF outperforms both EchoNeRF_LoS and NeRF2. Interestingly, MLP incurs a lower RPE than NeRF2  suggesting that RSSI is amenable to interpolation, and NeRF2’s implicit representation may not be an advantage for this interpolation task.
2000 receiver locations 1000 receiver locations
Method Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$ Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$
MLP - - 1.03 - - 0.65
Heatmap Seg. 0.12 $\pm$ 0.03 0.21 $\pm$ 0.05 1.32 0.09 $\pm$ 0.02 0.16 $\pm$ 0.04 1.46
NeRF2 0.14 $\pm$ 0.02 0.24 $\pm$ 0.03 4.36 0.12 $\pm$ 0.02 0.21 $\pm$ 0.04 4.2
EchoNeRF_LoS 0.27 $\pm$ 0.07 0.42 $\pm$ 0.10 9.12 0.25 $\pm$ 0.04 0.39 $\pm$ 0.06 10.86
EchoNeRF 0.38 $\pm$ 0.06 0.55 $\pm$ 0.06 3.56 0.32 $\pm$ 0.06 0.48 $\pm$ 0.05 4.32
Table 1: Performance Results for Wall_IoU, F1 Score, and RPE
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative comparison of ground truth floorplans against baselines. The bottom two rows show floorplans learnt by EchoNeRF_LoS and EchoNeRF with sharper walls and boundaries.
4.2 Qualitative Results: Visual floorplans, RSSI heatmap, and basic ray tracing
$\blacksquare$ Visual floorplans. Figure 5 presents visualization from all baselines and a comparison with our LoS-only model (as ablation). All the floorplans use $N=2000$ receiver locations. We make the following observations. (1) Heatmap Segmentation leverages the difference of RSSI on opposite sides of a wall, however, reflections pollute this pattern, especially at larger distances between $Tx$ and $Rx$. Further, signals leak through open doors, injecting errors in the room boundaries. (2) NeRF2 performs poorly since its MLP learns one among many possible assignments of virtual transmitters to fit the RSSI training data. The virtual transmitters hardly correlate to the walls of the environment. (3) EchoNeRF_LoS can infer the position of inner walls. However, these walls are thick and slanted because while EchoNeRF_LoS can identify occlusions between a $\langle Tx,Rx\rangle$ pair, it cannot tell the shape and pattern of these occlusions. Crucially, EchoNeRF_LoS also cannot infer the boundary walls since no receivers are located outside the house. (4) EchoNeRF outperforms the baselines, sharpens the inner walls compared to EchoNeRF_LoS, and constructs the boundary walls well.
Shortcomings: Recall that some parts of the floorplan are in the “blind spots” of our dataset since no reflection arrives from those parts to any of our sparse $Rx$ locations (e.g., see bottom left corner of the $1^{st}$ floorplan; no signals reflect off this region to arrive at any of the $Rx$ locations). Hence, EchoNeRF is unable to construct the bottom of the left wall in this floorplan. Finally, note that areas outside the floorplan (e.g., the regions on the right of $6^{th}$ floorplan) cannot be estimated correctly since no measurements are available from those regions (hence, those voxels do not influence the gradients).
$\blacksquare$ RSSI prediction. Figure 7 visualizes and compares predicted RSSI. The top row shows predictions at new $Rx$ locations with the $Tx$ held at the trained location; the bottom row shows predictions when both $Tx$ and $Rx$ are moved to new locations. Two key observations emerge: (1) EchoNeRF is limited by Sionna’s inability to simulate through-wall signal penetration; NeRF2 has access to an expensive license for a through-wall simulation and shows better predictions inside the rooms. However, in areas that EchoNeRF can "see" (e.g., corridors in the top row), the awareness of reflecting surfaces leads to significantly better predictions. (2) When the $Tx$ location differs from that used in training, EchoNeRF’s improvement over NeRF2 is significant. This is the core advantage of first solving the inverse problem and then leveraging it for the (forward) RSSI prediction.
Figure 6: Heatmaps highlighting EchoNeRF’s ability to learn signal propagation. (Top row) Inferred RSSI heatmaps with $Tx$ (red star) as used in training. (Bottom row) A new $Tx$ (green star) degrades NeRF2 and MLP while EchoNeRF shows accurate predictions.
Figure 7: (a) Tracing reflections on the learnt floorplan. (b) True reflections from Sionna.
$\blacksquare$ Learning reflected rays. For a given $\langle Tx,Rx\rangle$ pair, we examine the points in the plausible set $\mathcal{V}$ that contribute to the reflections. Fig. 7 compares the ray-tracing results from the NVIDIA Sionna simulator (we pick only first order reflections). EchoNeRF captures many of the correct reflections. Of course, some are incorrect – a false positive occurs in the bottom right room since some wall segment is missing in our estimate; false negatives also occur in the top right room where again some parts of the wall are missing.
4.3 Relaxing Assumptions & Sensitivity Study
$\blacksquare$ Transmitter’s location. We assumed knowledge of $Tx$ locations, however, we relax this by applying maximum likelihood estimation on observed RSSI power, $\psi^{*}$ (see Appendix D). On average, the estimated Tx location error is $2.08$ pixels in floorplans of sizes $\approx 512\times 512$ pixels.
Error $\sigma$(m) 0 0.5 1 2
Wall_IoU 0.38 0.35 0.33 0.29
Table 2: Estimated Wall_IoU at various levels of injected noise $\sigma$
Figure 8: EchoNeRF’s floorplan inference with furniture in conference (left) and apartment (right) layouts.
$\blacksquare$ Receiver location error. Table 2 shows EchoNeRF’s sensitivity to $Rx$ location errors. We inject Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ to the $Rx$ locations; $\sigma=1$ implies a physical error of $1$m. Wall_IoU accuracy obviously drops with error but $0.5$ meter of error is tolerable without destroying the floorplan structure. Advancements in WiFi positioning systems have demonstrated robust sub-meter error.
$\blacksquare$ Effect of Furniture. Fig. 8 visualizes inferred floorplans when toy objects are scattered in open spaces ($Rx$ locations remain $N$=$2000$). EchoNeRF is able to identify some of the object blobs but sharpening the small objects is challenging due to more higher order reflections from furniture. Follow up work is needed, either in modeling $2^{nd}$ order reflections or by imposing stronger regularizations.
5 Follow ups and Conclusion
Follow-ups. (1) The ability to model $2^{nd}$ order reflections will boost EchoNeRF’s accuracy, allowing it to sharpen the scene and decode smaller objects. For short range applications, such as non-intrusive medical imaging, $2^{nd}$ and $3^{rd}$ order reflections would be crucial. This remains an important direction for follow-on research. (2) Extending EchoNeRF to 3D floorplans is also of interest, and since it is undesirable to increase the number of measurements, effective 3D priors, or 2D-to-3D post-processing, may be necessary. Such post-processing tools exist cedreo2025 but we have not applied them since our goal is to improve NeRF’s inherent inverse solver. (3) Finally, EchoNeRF floorplans can offer valuable spatial context to Neural RIR synthesizers like zhao2023nerf2 ; newrf ; brunetto2024neraf ; AV-NeRF ; NACF . Synthesized RIR could in-turn aid EchoNeRF’s floorplan inference, forming the basis for an alternating optimization strategy. We leave these ideas to follow-up research.
Conclusion. In summary, we re-design the NeRF framework so it can learn to "see" its environment through line-of-sight (LoS) paths and reflections. This work takes a step towards neural wireless imaging with room for further research and varied downstream applications.
Appendix for
“Can NeRFs See without Cameras?”
Appendix A Background on Channel Impulse Response
When a wireless signal propagates, it is typically influenced by multipath effects such as reflections and scattering, as well as attenuation caused by the surrounding environment—collectively referred to as the channel. The overall impact of these phenomena on the signal is characterized by a linear model known as the Channel Impulse Response (CIR) [27].
Mathematically, the CIR is expressed as a sum of scaled and delayed impulses as shown in Eqn 9.
$h(t)=\sum_{i=1}^{N}a_{i}e^{j\phi_{i}}\delta(t-\tau_{i}),$ (9)
where $N$ is the number of multipath components, $a_{i}$ denotes the amplitude (attenuation factor) of the $i$-th path, $\phi_{i}$ represents the phase shift of the $i$-th path, and $\tau_{i}$ is the delay of the $i$-th path.
For an input signal $x(t)$ transmitted through the channel $h(t)$, the output signal measured at a Rx, $y(t)$ is obtained by the convolution:
$y(t)=x(t)*h(t)+w(t),$ (10)
where $w(t)$ represents zero-mean additive noise. For a simple two-path channel with a line of sight (LoS) path and one reflected path, the CIR $h(t)$ is given as
$h(t)=a_{1}\delta(t)+a_{2}e^{j\phi_{2}}\delta(t-\tau_{2}),$ (11)
The received signal $y(t)$ would then be:
$y(t)=x(t)*\left[a_{1}\delta(t)+a_{2}e^{j\phi_{2}}\delta(t-\tau_{2})\right]+w(t)$ (12)
Appendix B Modelling Wideband Multipath Signal Power
This section shows how received power in multipath scenarios can be approximated as a sum of powers of LoS and all other multipaths. In frequency domain, the received signal $y(t)$ at a particular receiver can be expressed as
$Y(f_{k})=H(f_{k})X(f_{k})+W(f_{k})$ (13)
Here, $X(f_{k})$ and $W(f_{k})$ represent the discrete Fourier transforms of the signal $x(t)$ and the additive noise $w(t)$ at subcarrier frequency $f_{k}$ with $k\in\{0,1,...,K\}$. The channel can be written as $H(f_{k})=\sum_{l=0}^{L}a_{lk}\exp(-j2\pi f_{k}\tau_{l})$ where $l\in\{1,2,...,L\}$ is an index over separate multipaths. Here, $a_{lk}$ and $\tau_{l}$ represents the attenuation and phase of the $l$-th multipath component at subcarrier $k$. We assume that the channel $H(f_{k})$ and the signal $X(f_{k})$ are independent.
The received power at each frequency $k$ is given by $\psi_{Y_{k}}=\operatorname{\mathbb{E}}[\left|Y(f_{k})\right|^{2}]$.
$\operatorname{\mathbb{E}}[\left|Y(f_{k})\right|^{2}]$ $=\operatorname{\mathbb{E}}[\left|H(f_{k})X(f_{k})+W(f_{k})\right|^{2}]$ (14)
$=\operatorname{\mathbb{E}}[\left|H(f_{k})\right|^{2}\left|X(f_{k})\right|^{2}]+\operatorname{\mathbb{E}}[\left|W(f_{k})\right|^{2}]+2\text{Re}\{\operatorname{\mathbb{E}}[H(f_{k})X(f_{k})W(f_{k})]\}$
$=\operatorname{\mathbb{E}}[\left|H(f_{k})\right|^{2}]\operatorname{\mathbb{E}}[\left|X(f_{k})\right|^{2}]+\operatorname{\mathbb{E}}[\left|W(f_{k})\right|^{2}]+2\text{Re}\{\operatorname{\mathbb{E}}[H(f_{k})X(f_{k})]\operatorname{\mathbb{E}}[W(f_{k})]\}$
$=\operatorname{\mathbb{E}}[\left|H(f_{k})\right|^{2}]\operatorname{\mathbb{E}}[\left|X(f_{k})\right|^{2}]+\operatorname{\mathbb{E}}[\left|W(f_{k})\right|^{2}],\because\operatorname{\mathbb{E}}[W(f_{k})]=0$
$=\psi_{X_{k}}\operatorname{\mathbb{E}}[\left|H(f_{k})\right|^{2}]+\psi_{W_{k}}$
where $\psi_{X_{k}}=\operatorname{\mathbb{E}}[\left|X(f_{k})\right|^{2}]$ and $\psi_{W_{k}}=\operatorname{\mathbb{E}}[\left|W(f_{k})\right|^{2}]$. Next,
[…]
$=\operatorname{\mathbb{E}}\left[\left(\sum_{l=0}^{L}a_{lk}\exp\left(-j2\pi f_{k}\tau_{l}\right)\right)\left(\sum_{m=0}^{L}a_{mk}^{*}\exp\left(j2\pi f_{k}\tau_{m}\right)\right)\right]$
$=\operatorname{\mathbb{E}}\left[\left(\sum_{l=0}^{L}\sum_{m=0}^{L}a_{lk}a_{mk}^{*}\exp\left(-j2\pi f_{k}\left(\tau_{l}-\tau_{m}\right)\right)\right)\right]$
$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}+\sum_{l=0}^{L}\sum_{\begin{subarray}{c}m=0\\ m\neq l\end{subarray}}^{L}a_{lk}a_{mk}^{*}\exp\left(-j2\pi f_{k}\left(\tau_{l}-\tau_{m}\right)\right)\right]$
$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]+\sum_{l=0}^{L}\sum_{\begin{subarray}{c}m=0\\ m\neq l\end{subarray}}^{L}\operatorname{\mathbb{E}}\left[a_{lk}a_{mk}^{*}\right]\exp\left(-j2\pi f_{k}\left(\tau_{l}-\tau_{m}\right)\right)$
$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]$
We used the fact that channel gains between different multipaths $l,m\ \text{with}\ l\neq m$ have vanishingly small correlation. Therefore, the total received power $\psi$ can be computed by summing all individual subcarrier powers.
$\psi$ $=\sum_{k=1}^{K}\psi_{Y_{k}}$
$=\sum_{k=1}^{K}\psi_{X_{k}}\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]+\psi_{W_{k}}$
$=\sum_{l=0}^{L}\sum_{k=1}^{K}\operatorname{\mathbb{E}}\left[\psi_{X_{k}}|a_{lk}|^{2}\right]+\psi_{W_{k}}$
Here, we separate out $\sum_{k=1}^{K}\operatorname{\mathbb{E}}\left[\psi_{X_{k}}|a_{0k}|^{2}\right]=\psi_{LoS}$, to represent the LoS power over all subcarriers and $\sum_{k=1}^{K}\operatorname{\mathbb{E}}\left[\psi_{X_{k}}|a_{lk}|^{2}\right]=\psi_{ref_{l}},1\leq l\leq L$, to denote the power of $l$-th order reflections.
Appendix C Approximating Channel with First-Order Reflections
EchoNeRF models the total received power at the Rx as the combination of the LoS power EchoNeRF_LoS, and the contributions from all the first-order reflections. To validate the contribution of the achievable power from EchoNeRF when compared to the total received power $\psi$, we evaluate the relative contributions of these signals to the total power using the NVIDIA Sionna simulator [15]. To this end, we compute the ratios of the LoS signal $\psi_{LoS}$, LoS with the first order reflections $\psi_{LoS}+\psi_{ref_{1}}$, and LoS with the first two orders of reflections $\psi_{LoS}+\psi_{ref_{1}}+\psi_{ref_{2}}$. These are compared to the total received power $\psi$, which is approximated as the sum of the LoS power and the power from the first ten reflections.
Fig 9 shows path power contribution ratio from different paths in histogram. While the $\psi_{LoS}$ power alone only accounts for approximately 70% of the total received power and is more spread out, $\psi_{LoS}+\psi_{ref_{1}}$ accounts to 95% of the total power, with a reduced spread. Moreover, secondary reflections $\psi_{ref_{2}}$ only contribute to less than 3% of the total power. Hence, EchoNeRF models the first-order reflections along with the line-of-sight.
Figure 9: Histograms illustrating the contribution ratios from line-of-sight (LoS), LoS combined with first-order reflections, and LoS combined with first- and second-order reflections. The orange graph highlights the significant contribution of first-order reflections to the total power, supporting EchoNeRF’s approach of modeling only the first reflection alongside the LoS power.
Appendix D Relaxing Tx Assumptions
We relax the assumption that Tx locations are known. Given the set of receiver locations $\{\mathbf{Rx}_{i}\}$ and the signal powers $\{\psi_{i}\}$, the goal is to estimate the transmitter location $\mathbf{Tx}=(\mathbf{Tx_{x}},\mathbf{Tx_{y}})$. To achieve this, we apply a maximum likelihood estimate (MLE). Briefly, among all the measured signal powers $\{\psi_{i}\}$ from a given Tx  we identify the $P$ strongest signal powers and their corresponding received locations. The rationale behind selecting the strongest powers is that they are significantly influenced by the LoS component, allowing us to model them effectively only using the Friss’ equation [1]. We assume independence among the measurements since the received LoS power across locations, for a given a Tx location, are independent. So, the likelihood equation for all these $P$ measurements can be written as:
$p(\psi_{1},\psi_{2},\dots,\psi_{P}|\mathbf{Tx})=\prod_{i=1}^{P}p(\psi_{i}|\mathbf{Tx})$
We approximate that the $\psi_{i}$ is normally distributed with a mean modeled by the line-of-sight power $\frac{K}{d_{i}^{2}}$ and variance $\sigma^{2}$ where $d_{i}=||\mathbf{Tx}-\mathbf{Rx}_{i}||$ is the distance between $\mathbf{Tx}$ and $\mathbf{Rx}_{i}$. The likelihood function for each observation $\psi_{i}$ is thus given by:
$p(\psi_{i}|\mathbf{Tx})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\psi_{i}-\frac{K}{d_{i}^{2}})^{2}}{2\sigma^{2}}\right)$
Maximizing log-likelihood $L$ of $\{\psi_{i}\}$ $\forall\ i\in\{1,\ldots,P\}$
$\log L(\mathbf{Tx})=\sum_{i=1}^{P}\log\left(\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\psi_{i}-\frac{K}{d_{i}^{2}})^{2}}{2\sigma^{2}}\right)\right)$
$\log L(\mathbf{Tx})=-\frac{P}{2}\log(2\pi\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{P}\left(\psi_{i}-\frac{K}{d_{i}^{2}}\right)^{2}$
Minimizing the second term gives the optimal $\mathbf{Tx}^{*}$ as:
${\mathbf{Tx}^{*}}=\operatorname*{argmin}_{\mathbf{Tx}}\sum_{i=1}^{P}\left(\psi_{i}-\frac{K}{\|\mathbf{Tx}-\mathbf{Rx}_{i}\|^{2}_{2}}\right)^{2}$
We use Scipy’s ‘minimize’ with the BFGS method to numerically solve for $\mathbf{Tx}^{*}$. Fig 13. visualizes the ground truth and the estimated Tx locations across 6 floorplans. The estimated Tx positions closely match the ground truth, and we report the Tx location error to be 2.08 pixels. Fig 10. demonstrates the performance of EchoNeRF_LoS and EchoNeRF using the estimated Tx locations for the 6 floorplans in Fig 13. Performance is comparable to that achieved with ground truth Tx locations, highlighting robustness.
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative comaprison of EchoNeRF_LoS and EchoNeRF when Tx locations are unknown, and are estimated. The top row shows the ground truth floorplan, Rx locations along with the estimated Txs in blue. The second and third row displays the performance of EchoNeRF_LoS and EchoNeRF respectively. Despite the Tx locations being unknown, our methods accurately estimate them, leading to performance comparable to the case where Tx locations are known.
Appendix E Details on Model Training
The signal power measured at the receiver is typically represented in a logarithmic scale. RSSI values generally range from -50 dB to -120 dB, where a higher value (e.g., -50 dB) corresponds to a stronger signal. Fig 11 illustrates a typical input to EchoNeRF  where measurements have been collected from approximately 2000 Rxs positioned in the floorplan, with data gathered from five Txs.
E.1 Linear-Scale RSSI Loss:
For the training of EchoNeRF, we optimize on the linear-scale RSSI values. Linear loss ensures that the receivers that capture stronger signals are given more importance during training. We partition our dataset into an 80-20 split, using 80% of the data for model training, including baselines.
For EchoNeRF’s network, we employ a simple 8-layer MLP with a hidden dimension of 256 units. For each voxel $v_{j}$, the outputs from the final layer are passed through a sigmoid activation to obtain the opacity $\delta$, and through a Gumbel softmax [16] layer to sample the output normal $\omega$ from one of the possible $K_{\omega}$ orientations. This sampled orientation is then used in the subsequent stages of training, such as for calculating the direction of the reflected signal $M$. For the learnable baselines, such as MLP and NeRF2, we adopt the same architecture as used in EchoNeRF.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Observed signal power at the Rxs . The top left figure shows the positions of the Rxs and Txs, followed by the power at the receivers from each of the five transmitters. The colormap ranges from red showing stronger signals to blue for weaker signals.
E.2 Supervising Voxel Orientations
EchoNeRF leverages the spatial gradient of a voxel’s opacity, $\nabla\delta$, to supervise its orientation during the multi-stage training process. To compute this gradient, we evaluate the opacities of neighboring voxels along each of the $K_{\omega}$ directions and apply finite difference methods. We found that this approach yielded superior results compared to using the gradient available via autograd.
Figure 12: An incoming ray from a transmitter Tx reflecting around voxel $v_{j}$ and arriving at receiver Rx. The incoming ray makes an incident angle $\theta$ with the normal $\omega_{j}$ to the reflecting surface. The ray after reflection passes a receiver Rx at a certain distance making an angle $\beta$.
In general, the power from reflections depends not only on the total distance traveled but also on the angle at which the reflection occurs at the voxel $v_{j}$, and whether the reflected ray reaches the receiver (Rx). We model this behavior through $\theta$ and $\beta$ respectively which parameterize a nonlinear function $f$. The incidence angle $\theta$ measures the angle between the Tx and the orientation $\omega_{j}$, and $\beta$ denotes the angle Rx makes with the reflected ray $M$ (see fig 12). Note that if $v_{j}\in\mathcal{V}$, and if $\omega_{j}$ is correct, $\beta=0$. The reflected ray $M$ can be computed as shown in Eqn 15.
$M=(v_{j}-Tx)-2\left[(v_{j}-Tx).\omega_{j}\right]\omega_{j}$ (15)
Here $\omega_{j}$ is a unit vector.
Figure 13: Comparison of Ground truth Tx locations indicated in red in the first column with the estimated Tx locations shown in blue from starting from column two. The Rx positions used for the estimation are marked in green.
E.3 Detailed network parameters
We assume the floorplan is an unknown shape inside a $512\times 512$ grid. For any ray or ray segment, we uniformly sample $n_{r}=64$ voxels on it. We choose $\lambda_{1}=\lambda_{2}=0.01$ for LoS training followed by $\lambda_{1}=\lambda_{2}=0.1$ for training the EchoNeRF model. We find that discretized opacity values to $\{0,1\}$ improve our LoS model. We use the straight-through estimator [4] to avoid the unavailability of the gradient at the discretization step. To help optimization and to encourage sparsity of the number of reflections, we use only the top-$k$ contributions ($k$=$10$) while training the reflection model. We use the ADAM optimizer [17] with $1.0^{-4}$ learning rate. We train our models on NVIDIA A100 GPUs.
E.4 Heatmap Segmentation Implementation Details
The raw trajectory signal power values are first interpolated to obtain a heatmap that provides a smoother representation of the input measurements. Of course, interpolating for regions without any data can lead to incorrect results, especially in larger unseen areas. A rule-based classifier is then applied for segmentation, using two criteria: (1) RSSI values above a threshold to identify potential room areas, and (2) smoothness of the RSSI signal, assessed through the second-order derivative, to ensure continuity within rooms. The initial segmentation is refined using morphological operations (via dilation and erosion) with a 3x3 kernel to smooth rough edges and eliminate small components. Overlapping regions are resolved by comparing gradient magnitudes, followed by additional morphological processing and connected component analysis to obtain the final, refined segmentation.
Code: We plan to release our code, data, and baselines upon acceptance of the paper to ensure reproducibility. In the meantime, Section E provides sufficient details to allow readers to reproduce our results, especially since our network components are simple MLPs.
Appendix F Evaluation on Additional Floorplans
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Qualitative comparison of Ground Truth floorplans against those inferred by baselines Heatmap Segmentation and NeRF2. We note that while NeRF2 is unable to predcit any reasonable floorplan, Heatmap Segmentationś shape is limited by the (convex hull of the) trajectory data (see bottom left of the second row, first column). Additionally, it fails to capture critical details, such as door openings. The 4th and 5th rows show floorplans by our proposed models EchoNeRF_LoS and EchoNeRF. EchoNeRF_LoS captures the rough shape of the floorplan, especially the interior walls, while EchoNeRF further improves these walls by adjusting their thickness and accurately correcting their shape. EchoNeRF also correctly identifies the floorplan boundary, as evidenced in the last column, where the exact boundary is captured just from the reflections. To understand the signal propagation captured by EchoNeRF, we place one Tx in each floor plan randomly (that is not present in the training data) and evaluate the signal power at discrete receivers. These Rxs are placed on a 2D grid at equal intervals and the predicted signal power is converted into a heatmap. The bottom row shows these inferred signal power heatmaps with the brightest point indicating the Tx location (as the Rx closest to the Tx receives the highest power). EchoNeRF is not only able to predict the signals well across the floorplan, but also capture the propagation paths i.e., LoS signal and the first-order reflections. For instance, in the first column, the left portion of the center hall receives power only due to the wall reflection from the left wall.
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Additional qualitative comparisons of Ground Truth floorplans against those inferred by baselines Heatmap Segmentation and NeRF2. The 4th and 5th rows show floorplans by our proposed models EchoNeRF_LoS and EchoNeRF with clearly identified walls and boundaries. The bottom row shows inferred signal power heatmaps demonstrating EchoNeRF’s capability to learn accurate signal propagation.
Appendix G Societal Impact Statement
We acknowledge that NeRFs hold significant potential for positive societal impact. Applications span AR/VR, medical imaging, airport security, and education, where accurate 3D reconstructions can greatly enhance functionality and understanding. However, our work on EchoNeRF also introduces potential risks. In particular, the ability to infer detailed spatial layouts from limited sensory input could be misused to access private or sensitive floorplan information. We emphasize the importance of responsible use.

### v2

@@
-A Posterior Sampling Approach to Inferring Indoor Floorplans from Walk Trajectories
+Can NeRFs See without Cameras?
 Abstract
@@
 Figure 1: LoS and correct multipath reflections.
-Figure 2: (a) Fitting signal power using virtual $Tx$s.
-(b) Ideally, virtual $Tx$s should be located on wall surfaces.
-(c) Sparse virtual $Tx$s learnt by NeWRF shown in black/gray dots.
-(d) Dense virtual $Tx$s learnt by NeRF2.
+Figure 2: (a) Fitting signal power using virtual TxTxs.
+(b) Ideally, virtual TxTxs should be located on wall surfaces.
+(c) Sparse virtual TxTxs learnt by NeWRF shown in black/gray dots.
+(d) Dense virtual TxTxs learnt by NeRF2.
 Neither correspond to the true (red) floorplan.
@@
 This function only models the LoS and the first order reflections.
-We disregard the higher orders since they are very complex to model and contribute, on average, $<6\%$ of the total power (see statistics in Appendix C).
+We disregard the higher orders since they are very complex to model and contribute, on average, $6\%$ of the total power (see statistics in Appendix C).
 The NeRF model we use is a remarkably simple MLP designed to predict the density $\delta\in[0,1]$ and orientation $\omega\in[-\pi,\pi]$ of a specified voxel in the indoor scene.
@@
 4).
-This converges quickly because the network easily learns the transparent voxels ($\delta$=00) located along LoS paths.
+This converges quickly because the network easily learns the transparent voxels ($\delta$=0) located along LoS paths.
 For LoS paths that are occluded, the network incorrectly learns excessive opaque voxels between the $\langle Tx,Rx\rangle$, but this does not affect the LoS error since the path is anyway occluded.
@@
 We randomly place $M$ Txs, one in each room, denoted as $T_{m}$.
-To simulate omnidirectional transmissions at $2.4$GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan.
+To simulate omnidirectional transmissions at 2.42.4GHz from each Tx location, we shoot $10^{7}$ rays into the given floorplan.
 For receiver locations, we sample the user trajectory at a fixed time interval to obtain $N$ Rx locations, denoted as $R_{n}$.
@@
 The following equation defines the metric:
-[…]
+$\text{{Wall\_IoU}}=\frac{WP\cap WP^{*}}{WP\cup WP^{*}}$
 where $WP$ denotes the set of predicted wall pixels and $WP^{*}$ denotes the true wall pixels.
@@
 Figure 6: Heatmaps highlighting EchoNeRF’s ability to learn signal propagation.
-(Top row) Inferred RSSI heatmaps with $Tx$ (red star) as used in training.
-(Bottom row) A new $Tx$ (green star) degrades NeRF2 and MLP while EchoNeRF shows accurate predictions.
+(Top row) Inferred RSSI heatmaps with TxTx (red star) as used in training.
+(Bottom row) A new TxTx (green star) degrades NeRF2 and MLP while EchoNeRF shows accurate predictions.
 Figure 7: (a) Tracing reflections on the learnt floorplan. (b) True reflections from Sionna.
@@
 Table 2 shows EchoNeRF’s sensitivity to $Rx$ location errors.
-We inject Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ to the $Rx$ locations; $\sigma=1$ implies a physical error of $1$m.
+We inject Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ to the $Rx$ locations; $\sigma=1$ implies a physical error of 11m.
 Wall_IoU accuracy obviously drops with error but $0.5$ meter of error is tolerable without destroying the floorplan structure.
@@
 Fig.
-8 visualizes inferred floorplans when toy objects are scattered in open spaces ($Rx$ locations remain $N$=$2000$).
+8 visualizes inferred floorplans when toy objects are scattered in open spaces ($Rx$ locations remain NN=$2000$).
 EchoNeRF is able to identify some of the object blobs but sharpening the small objects is challenging due to more higher order reflections from furniture.
@@
 $=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]+\sum_{l=0}^{L}\sum_{\begin{subarray}{c}m=0\\ m\neq l\end{subarray}}^{L}\operatorname{\mathbb{E}}\left[a_{lk}a_{mk}^{*}\right]\exp\left(-j2\pi f_{k}\left(\tau_{l}-\tau_{m}\right)\right)$
-$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]$
-We used the fact that channel gains between different multipaths $l,m\ \text{with}\ l\neq m$ have vanishingly small correlation.
-Therefore, the total received power $\psi$ can be computed by summing all individual subcarrier powers.
+$=\operatorname{\mathbb{E}}\left[\sum_{l=0}^{L}|a_{lk}|^{2}\right]\$
+[…]
 $\psi$ $=\sum_{k=1}^{K}\psi_{Y_{k}}$
@@
 $p(\psi_{1},\psi_{2},\dots,\psi_{P}|\mathbf{Tx})=\prod_{i=1}^{P}p(\psi_{i}|\mathbf{Tx})$
-We approximate that the $\psi_{i}$ is normally distributed with a mean modeled by the line-of-sight power $\frac{K}{d_{i}^{2}}$ and variance $\sigma^{2}$ where $d_{i}=||\mathbf{Tx}-\mathbf{Rx}_{i}||$ is the distance between $\mathbf{Tx}$ and $\mathbf{Rx}_{i}$.
-The likelihood function for each observation $\psi_{i}$ is thus given by:
+[…]
 $p(\psi_{i}|\mathbf{Tx})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\psi_{i}-\frac{K}{d_{i}^{2}})^{2}}{2\sigma^{2}}\right)$
@@
 We use the straight-through estimator [4] to avoid the unavailability of the gradient at the discretization step.
-To help optimization and to encourage sparsity of the number of reflections, we use only the top-$k$ contributions ($k$=$10$) while training the reflection model.
+To help optimization and to encourage sparsity of the number of reflections, we use only the top-$k$ contributions (kk=$10$) while training the reflection model.
 We use the ADAM optimizer [17] with $1.0^{-4}$ learning rate.

### v3

@@
-Can NeRFs See without Cameras?
+Can NeRFs “See” without Cameras?
 Abstract
@@
 1 Introduction
-NeRFs mildenhall2021nerf ; nerfInvProb ; mipnerf ; pixelnerf have delivered impressive results in solving inverse problems, resulting in 3D scene rendering.
+NeRFs Mildenhall et al.
+(2021); Wang et al.
+(2023); Barron et al.
+(2021); Yu et al.
+(2021) have delivered impressive results in solving inverse problems, resulting in 3D scene rendering.
 While NeRFs have mostly used pictures (from cameras or LIDARs) to infer a 3D scene, we ask if the core ideas can generalize to the case of wireless signals (such as RF or audio).
@@
 Using these $N$ wireless measurements, is it possible to learn a representation of the scene, such as the floorplan of the user’s home?
-An optical NeRF understandably fails since it is not equipped to handle multipath.
+A vanilla NeRF understandably fails since it is not equipped to handle multipath.
 This paper is focused on redesigning NeRFs so they can learn to image the environment, thereby solving the inverse problem from ambient wireless signals.
-A growing body of work specnerf ; winert ; zhao2023nerf2 ; newrf ; luo2022learning ; AV-NeRF is investigating connections between NeRFs and wireless.
-While none have concentrated on imaging, NeRF2 zhao2023nerf2 and NeWRF newrf have augmented NeRFs to correctly synthesize WiFi signals at different locations inside an indoor space.
+A growing body of work Basu and Chakraborty (2024); Orekondy et al.
+(2023); Zhao et al.
+(2023); Lu et al.
+(2024); Luo et al.
+(2022); Liang et al.
+(2023a) is investigating connections between NeRFs and wireless.
+While none have concentrated on imaging, NeRF2 Zhao et al.
+(2023) and NeWRF Lu et al.
+(2024) have augmented NeRFs to correctly synthesize WiFi signals at different locations inside an indoor space.
 However, correct synthesis is possible without necessarily learning the correct signal propagation models.
@@
 Lastly, to handle some gradient imbalance issues, EchoNeRF freezes the LoS model once it converges, and uses this intermediate state to partly supervise the reflection model.
-To evaluate EchoNeRF, we train on $2.4$ GHz WiFi signals from NVIDIA’s Sionna simulator hoydis2023sionna , with floorplans from the Zillow’s Indoor Dataset (ZIND) cruz2021zillow .
+To evaluate EchoNeRF, we train on $2.4$ GHz WiFi signals from NVIDIA’s Sionna simulator Hoydis et al.
+(2023), with floorplans from the Zillow’s Indoor Dataset (ZIND) Cruz et al.
+(2021).
 Results show consistent improvement over baselines in terms of the estimated floorplan’s IoU and F1 score.
@@
 Wireless (WiFi) channel prediction using NeRFs.
-NeWRFnewrf and NeRF2zhao2023nerf2 are recent papers that have used NeRFs to predict the wireless channel impulse response (CIR) Tse_Viswanath_2005 at unknown locations inside a room.
+NeWRFLu et al.
+(2024) and NeRF2Zhao et al.
+(2023) are recent papers that have used NeRFs to predict the wireless channel impulse response (CIR) Tse and Viswanath (2005) at unknown locations inside a room.
 Drawing a parallel to optical NeRFs, a voxel’s color in optics becomes a voxel’s transmit power in wireless.
@@
 Neural radiance fields for audio.
-Another active line of research focuses on predicting room impulse response (RIR) for audio luo2022learning ; INRAS ; AV-NeRF ; brunetto2024neraf .
-Neural Acoustic Field (NAF) luo2022learning extended the classical NeRF to train on RIR measurements in a room and predict the RIR (magnitude and phase) at new $\langle Tx,Rx\rangle$ locations.
+Another active line of research focuses on predicting room impulse response (RIR) for audio Luo et al.
+(2022); Su et al.
+(2022); Liang et al.
+(2023a); Brunetto et al.
+(2024).
+Neural Acoustic Field (NAF) Luo et al.
+(2022) extended the classical NeRF to train on RIR measurements in a room and predict the RIR (magnitude and phase) at new $\langle Tx,Rx\rangle$ locations.
 NAF identified the possibility of overfitting to the RIR and proposed to learn, jointly, the local geometric features of the environment (as spatial latents) and the NAF parameters.
@@
 EchoNeRF requires no floorplan supervision, and secondly, relies entirely on signal power (less informative than RIR) to estimate the floorplan.
-Follow up work are embracing more information about the surroundings (pictures NACF ; INRAS ; AV-NeRF ; fewshotvisual , LIDAR scans oufqir2020arkit , meshes and optical NeRFs brunetto2024neraf ), to boost RIR accuracy.
+Follow up work are embracing more information about the surroundings (pictures Liang et al.
+(2023b); Su et al.
+(2022); Liang et al.
+(2023a); Majumder et al.
+(2022), LIDAR scans Oufqir et al.
+(2020), meshes and optical NeRFs Brunetto et al.
+(2024)), to boost RIR accuracy.
 Results are steadily improving, however, this sequence of ideas is unaligned with solving the core inverse problem.
 Our goal is to first invert the signal power to a floorplan, which can then enable CIR/RIR predictions.
-Modeling reflections in optical NeRFs Optical NeRFs have tackled reflections refneus ; nerfds ; flipnerf for synthesizing glossy surfaces and mirrors, and for re-lighting rudnev2022nerf ; relight .
-NeRFRen nerfren proposes to decompose a viewed image into a transmitted and a reflected component.
-Ref-NeRF verbin2022refnerf also focuses on reflections through a similar decomposition of the transmitted and reflected color, however, the reflected color is modeled as a function of the viewing angle and the surface-normal, resulting in accurate models of specular reflection.
-Several other papers have developed similar ideas boss2021nerd ; zhang2021nerfactor and the core insight centers around solving a two-component decomposition problem.
+Modeling reflections in optical NeRFs Optical NeRFs have tackled reflections Ge et al.
+(2023); Yan et al.
+(2023); Seo et al.
+(2023) for synthesizing glossy surfaces and mirrors, and for re-lighting Rudnev et al.
+(2022); Toschi et al.
+(2023).
+NeRFRen Guo et al.
+(2022b) proposes to decompose a viewed image into a transmitted and a reflected component.
+Ref-NeRF Verbin et al.
+(2022) also focuses on reflections through a similar decomposition of the transmitted and reflected color, however, the reflected color is modeled as a function of the viewing angle and the surface-normal, resulting in accurate models of specular reflection.
+Recent proposals such as NeRF-Casting Verbin et al.
+(2024) and oRCA Tiwary et al.
+(2023) have further improved the models of multipath for glossy and mirrored surfaces, and others Boss et al.
+(2021); Zhang et al.
+(2021) have developed similar ideas, with the core insights centering around solving a two-component decomposition problem.
 EchoNeRF faces the challenge of not knowing the number of rays adding up from all possible directions in the environment.
 Hence, EchoNeRF must solve a many-component decomposition problem by leveraging the physics of multipath signal propagation.
+In the context of LIDARs, PlatoNeRF Klinghoffer et al.
+(2024) and NeTF Shen et al.
+(2021) cope with unknown reflections, however, LIDARs have very high time resolution (due to high clock frequency), and is therefore able to assign the incoming rays to different time buckets.
+This temporal separation allows the NeRF to make separate measurements for different surfaces in the scene.
+Since EchoNeRF uses only signal power—a single scalar measurement that contains a mixture of the LoS and all the reflections—the inverse problem of recovering the voxel attributes from only the power is far more complex.
 3 EchoNeRF Model
@@
 $\psi=\psi_{LoS}+\psi_{ref_{1}}+\cdots+\psi_{ref_{n}}$
-where $\psi_{LoS}$ is the power from the direct line-of-sight (LoS) path, and $\psi_{ref_{k}}$ is the aggregate power from all $k^{th}$ order reflections (i.e., all signal paths that underwent exactly $k$ reflections before arriving at the $Rx$)  [^1] This model is an approximation since it relies on averaging measurements as explained in Appendix B.
-However, for this sensing application involving moving receivers, and when using wideband-multipath systems such as WiFi, this approximation holds well.
+where $\psi_{LoS}$ is the power from the direct line-of-sight (LoS) path, and $\psi_{ref_{k}}$ is the aggregate power from all $k^{th}$ order reflections (i.e., all signal paths that underwent exactly $k$ reflections before arriving at the $Rx$)  [^1] This model is a simplification since it ignores the signal phase in estimating the received power.
+Appendix B shows that with a moving wide-band receiver, like WiFi, the approximation may be tolerable for a sensing application like EchoNeRF.
 We assume $M$ fixed transmitters and move the $Rx$ to $N$ known locations and measure $\psi$ at each of them.
@@
 3.1 The LoS Model
-Friss’ equation antennaBook from electromagnetics models the free-space received power as $P_{r}=\frac{K}{d^{2}}$ where $d$ is the distance of signal propagation, and $K$ is a product of transmit-power, wavelength, and antenna-related constants antennaBook .
+Friss’ equation Balanis (2016) from electromagnetics models the free-space received power as $P_{r}=\frac{K}{d^{2}}$ where $d$ is the distance of signal propagation, and $K$ is a product of transmit-power, wavelength, and antenna-related constants Balanis (2016).
 We model this free-space (LoS) behavior in the NeRF framework through the following equation.
@@
 This makes the optimization problem excessively under-determined.
-We address this by modeling $\omega_{j}$ as a discrete value, i.e., $\omega_{j}\in\{1,2\dots K_{\omega}\}$.
-Larger $K_{\omega}$ is needed when the environment has complex surface orientations; however, most floorplans exhibit perpendicular walls manhattan1 ; manhattan2 ; manhattan3 and $K_{\omega}=4$ is adequate.
+We address this by modeling $\omega_{j}$ as a discrete value—multiples of $\frac{\pi}{K_{\omega}}$.
+Larger $K_{\omega}$ is needed when the environment has complex surface orientations; however, most floorplans exhibit perpendicular walls Zou et al.
+(2018); Coughlan and Yuille (1999); Guo et al.
+(2022a) and $K_{\omega}=4$ is adequate.
 Once $\omega_{j}$ becomes discrete, the voxels that can produce plausible reflections become far fewer – we call this the “plausible set” $\mathcal{V}$.
 Fig.
-4 visualizes $\mathcal{V}$ and shows $3$ out of many plausible reflections from voxels with $\omega_{j}=45^{\circ}$.
+4 visualizes $\mathcal{V}$ and shows $3$ out of many plausible reflections from voxels with $\omega_{j}=135^{\circ}$.
 Eq.
@@
 $\blacksquare$ Regularization: Floorplans demonstrate significant local similarity in orientation, hence we penalize differences in orientation among neighbors, using a regularization (Eq.
-8) similar to Total Variation TV_reg .
+8) similar to Total Variation Rudin et al.
+(1992).
 This can be achieved without additional computational cost to the neural network by directly utilizing voxel orientations obtained from each ray.
@@
 Floorplan and Wireless Simulation Dataset.
-Floorplans are drawn from the Zillow Indoor Dataset  cruz2021zillow .
-In each floorplan, we use the $A^{*}$ algorithm  astar to generate a walking trajectory that traverses all rooms.
-We use the NVIDIA Sionna RT hoydis2022sionna ; hoydis2023sionna – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)).
+Floorplans are drawn from the Zillow Indoor Dataset  Cruz et al.
+(2021).
+In each floorplan, we use the $A^{*}$ algorithm  Hart et al.
+(1968) to generate a walking trajectory that traverses all rooms.
+We use the NVIDIA Sionna RT Hoydis et al.
+(2022, 2023) – a ray tracer for radio propagation modeling – to compute the ground truth signal power (also known as received signal strength index (RSSI)).
 We randomly place $M$ Txs, one in each room, denoted as $T_{m}$.
@@
   1. 1.
-    NeRF2 zhao2023nerf2 : Models WiFi reflections via virtual transmitters to predict channel impulse response (CIR).
+    NeRF2 Zhao et al. (2023): Models WiFi reflections via virtual transmitters to predict channel impulse response (CIR).
   2. 2.
-    Heatmap Segmentation liu2012review : Interpolates CIR across the whole floorplan and applies an image segmentation algorithm (on the interpolated RSSI heatmap) to isolate each room.
+    Heatmap Segmentation Liu et al.
+(2012): Interpolates CIR across the whole floorplan and applies an image segmentation algorithm (on the interpolated RSSI heatmap) to isolate each room.
 Essentially, the algorithm identifies the contours of sharp RSSI change since such contours are likely to correspond to walls.
@@
 This is a harsh metric given wall pixels are a small fraction of the total floorplan; if a predicted wall is even offset by one pixel from the true wall, the Wall_IoU drops significantly.
-IoU iou has often been defined in terms of room pixels (instead of wall pixels); this is an overestimate in our opinion, since predicting even an empty floorplan results in an impressively high IoU.
-(B) F1 score metrics : Defined as $F1=\frac{2\times P\times R}{P+R}$, where $P$ is the precision and $R$ is the recall of the bitmap. $P$ and $R$ are defined based on wall pixels, similar as above.
+IoU Rezatofighi et al.
+(2019) has often been defined in terms of room pixels (instead of wall pixels); this is an overestimate in our opinion, since predicting even an empty floorplan results in an impressively high IoU.
+(B) F1 score Sokolova and Lapalme (2009): Defined as $F1=\frac{2\times P\times R}{P+R}$, where $P$ is the precision and $R$ is the recall of the bitmap.
+$P$ and $R$ are defined based on wall pixels, similar as above.
 (C) RSSI Prediction Error (RPE): We split all $Rx$ locations into a training and test set. RPE reports the average median RSSI error over all the test locations across floorplans.
 4.1 Overall Summarized Results
+2000 receiver locations 1000 receiver locations
+Method Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$ Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$
+MLP - - 1.03 - - 0.65
+Heatmap Seg. 0.12 $\pm$ 0.03 0.21 $\pm$ 0.05 1.32 0.09 $\pm$ 0.02 0.16 $\pm$ 0.04 1.46
+NeRF2 0.14 $\pm$ 0.02 0.24 $\pm$ 0.03 4.36 0.12 $\pm$ 0.02 0.21 $\pm$ 0.04 4.2
+EchoNeRF_LoS 0.27 $\pm$ 0.07 0.42 $\pm$ 0.10 9.12 0.25 $\pm$ 0.04 0.39 $\pm$ 0.06 10.86
+EchoNeRF 0.38 $\pm$ 0.06 0.55 $\pm$ 0.06 3.56 0.32 $\pm$ 0.06 0.48 $\pm$ 0.05 4.32
+Table 1: Performance Results for Wall_IoU, F1 Score, and RPE
 Table 1 reports comparative results between EchoNeRF and baselines, averaged over $20$ different experiments, using all $3$ metrics.
@@
 Interestingly, MLP incurs a lower RPE than NeRF2  suggesting that RSSI is amenable to interpolation, and NeRF2’s implicit representation may not be an advantage for this interpolation task.
-2000 receiver locations 1000 receiver locations
-Method Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$ Wall_IoU $\uparrow$ F1 Score $\uparrow$ RPE $\downarrow$
-MLP - - 1.03 - - 0.65
-Heatmap Seg. 0.12 $\pm$ 0.03 0.21 $\pm$ 0.05 1.32 0.09 $\pm$ 0.02 0.16 $\pm$ 0.04 1.46
-NeRF2 0.14 $\pm$ 0.02 0.24 $\pm$ 0.03 4.36 0.12 $\pm$ 0.02 0.21 $\pm$ 0.04 4.2
-EchoNeRF_LoS 0.27 $\pm$ 0.07 0.42 $\pm$ 0.10 9.12 0.25 $\pm$ 0.04 0.39 $\pm$ 0.06 10.86
-EchoNeRF 0.38 $\pm$ 0.06 0.55 $\pm$ 0.06 3.56 0.32 $\pm$ 0.06 0.48 $\pm$ 0.05 4.32
-Table 1: Performance Results for Wall_IoU, F1 Score, and RPE
 Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Qualitative comparison of ground truth floorplans against baselines. The bottom two rows show floorplans learnt by EchoNeRF_LoS and EchoNeRF with sharper walls and boundaries.
+6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 5: Qualitative comparison of ground truth floorplans against baselines.
+In the first row, red stars denote Tx locations and light gray dots denote Rx measurement locations.
+The bottom two rows show floorplans learnt by EchoNeRF_LoS (i.e., Stage 1) and EchoNeRF (i.e., Stage 2) with sharper walls and boundaries.
+More visualizations available at https://echonerf.github.io/
 4.2 Qualitative Results: Visual floorplans, RSSI heatmap, and basic ray tracing
@@
 Follow up work is needed, either in modeling $2^{nd}$ order reflections or by imposing stronger regularizations.
+$\blacksquare$ Material Sensitivity.
+Table 4 shows the mean Wall_IoU of EchoNeRF on five different materials averaged across six scenes (shown in Fig.
+5) Materials with higher reflectivity, such as concrete and glass, yield better performance than absorptive materials like wood.
+This is because more reflections allow better performance for EchoNeRFś reflection model.
+$\blacksquare$ Robustness to RSSI error.
+We added Gaussian noise to the RSSI measurements with a mean equal to the noise floor (in dB) and a variance of $4$ dB.
+We vary the noise floor levels ranging from 80-80 dB to 130-130 dB across the 6 floorplans shown in Fig.
+5.
+The SNR at a receiver is computed as the difference between the received signal power and the noise floor (e.g., a received power of 70-70 dB with a noise floor of 80-80 dB results in an SNR of $10$ dB).
+We report the mean Wall_IoU for EchoNeRF_LoS and EchoNeRF in Table 4.
+The performance drops with decreasing SNR; both EchoNeRF_LoS and EchoNeRF’s floorplans are still legible till $30$ dB, but below that (when noise power becomes comparable to the signal power) they break down, leading to missing and illegible walls.
+Our results are conservative; when we report a specific SNR level (e.g., 40dB), it represents the highest SNR (best-case) among all receivers in that scenario, meaning other receivers experience even lower SNRs.
+In practice, WiFi SNR ranges from 30-60dB depending on the distance from the router, with close-proximity measurements often exceeding 50-60dB.
+For average real-world SNR conditions around 45dB, the corresponding best-case SNR would be 60+dB, which aligns with the top rows where EchoNeRF demonstrates strong performance.
+[…]
+Concrete 0.251 0.371
+Glass 0.236 0.364
+Brick 0.232 0.357
+Marble 0.226 0.328
+Wood 0.227 0.311
+Table 3: Sensitivity across materials.
+[…]
+$\infty$ (no noise) 0.251 0.371 Legible
+60 0.246 0.336 Legible
+50 0.231 0.292 Legible
+40 0.226 0.298 Legible
+30 0.207 0.241 Missing walls
+10 0.090 0.140 Illegible
+Table 4: Noise robustness across SNR levels.
 5 Follow ups and Conclusion
@@
 (2) Extending EchoNeRF to 3D floorplans is also of interest, and since it is undesirable to increase the number of measurements, effective 3D priors, or 2D-to-3D post-processing, may be necessary.
-Such post-processing tools exist cedreo2025 but we have not applied them since our goal is to improve NeRF’s inherent inverse solver.
-(3) Finally, EchoNeRF floorplans can offer valuable spatial context to Neural RIR synthesizers like zhao2023nerf2 ; newrf ; brunetto2024neraf ; AV-NeRF ; NACF .
+Such post-processing tools exist Cedreo (2025) but we have not applied them since our goal is to improve NeRF’s inherent inverse solver.
+(3) Expanding evaluations beyond ZInD, which contains largely unfurnished, rectangular rooms, to richer datasets such as HM3D Ramakrishnan et al.
+(2021) and MVL Solarte et al.
+(2024) would also be valuable: furniture and clutter can introduce significant multipath effects that complicate RF signal modeling, and additional research is needed to understand these effects.
+We defer this to future work, as our focus in this paper is not so much to understand the limits of RF-based NeRFs, but to establish the feasibility of such frameworks.
+(4) Finally, EchoNeRF floorplans can offer valuable spatial context to Neural RIR synthesizers like Zhao et al.
+(2023); Lu et al.
+(2024); Brunetto et al.
+(2024); Liang et al.
+(2023a, b).
 Synthesized RIR could in-turn aid EchoNeRF’s floorplan inference, forming the basis for an alternating optimization strategy.
@@
 Conclusion.
-In summary, we re-design the NeRF framework so it can learn to "see" its environment through line-of-sight (LoS) paths and reflections.
-This work takes a step towards neural wireless imaging with room for further research and varied downstream applications.
+In summary, we re-design the NeRF framework so it can learn to "see" its environment by leveraging both line-of-sight (LoS) paths and multipath reflections.
+While such reflections bring to the sensor more information about the surroundings, their mixtures with the LoS path also complicates the core inverse problems.
+EchoNeRF takes a step towards solving this inverse problem, but also leaves room for further research in neural wireless imaging and varies downstream applications.
+Acknowledgments
+We would like to thank Prof.
+Shenlong Wang for his guidance in the initial phase of this work.
+We also thank the anonymous reviewers for their valuable feedback.
+This work was partially supported by NSF #2008338, #1909568, #2148583, and #MRI-2018966.
+This work used DELTA at NCSA through allocation CIS230230 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S.
+National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296
 Appendix for
-“Can NeRFs See without Cameras?”
+Can NeRFs “See” without Cameras?
 Appendix A Background on Channel Impulse Response
@@
 EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 10: Qualitative comaprison of EchoNeRF_LoS and EchoNeRF when Tx locations are unknown, and are estimated.
@@
 Overlapping regions are resolved by comparing gradient magnitudes, followed by additional morphological processing and connected component analysis to obtain the final, refined segmentation.
-Code: We plan to release our code, data, and baselines upon acceptance of the paper to ensure reproducibility.
+Code: We plan to release our code, data, and baselines soon.
 In the meantime, Section E provides sufficient details to allow readers to reproduce our results, especially since our network components are simple MLPs.
@@
 Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 14: Qualitative comparison of Ground Truth floorplans against those inferred by baselines Heatmap Segmentation and NeRF2.
@@
 Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. Heatmap Seg. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. NeRF2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 EchoNeRF LoS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. EchoNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6cm. EchoNeRF Signal Prediction Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Figure 15: Additional qualitative comparisons of Ground Truth floorplans against those inferred by baselines Heatmap Segmentation and NeRF2.